OpenStudy (anonymous):

Express in terms of i. 4sqrt-225

OpenStudy (anonymous):

\[4\sqrt{-225}\]

OpenStudy (anonymous):

how is this done

OpenStudy (anonymous):

okay, here's how to deal with square root of negative numbers \[\sqrt{-b}=i \sqrt{b}\]

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

try it on your expression

OpenStudy (anonymous):

the sqrt of -225 is 15

OpenStudy (anonymous):

15 and ?

OpenStudy (anonymous):

15

OpenStudy (anonymous):

and 5

OpenStudy (anonymous):

sqrt[225]=15 what's sqrt[-225]= ? look at my example above

OpenStudy (anonymous):

you are one of the few that actually help the person understand the problem

OpenStudy (anonymous):

okay so would it be sqrt-225 and 4 ?

OpenStudy (anonymous):

4 sqrt and 15 right

OpenStudy (anonymous):

\[\sqrt{-b}=i \sqrt{b}\] \[\sqrt{-225}=i\sqrt{225}=15i\]

OpenStudy (anonymous):

okay now I have the understanding of it I was trying to think of what you was saying my bad

OpenStudy (anonymous):

Here's more elaboration of what I did \[\sqrt{-1}=i\] \[\sqrt{-225}=\sqrt{225}*\sqrt{-1}=\sqrt{225}i=15i\]

OpenStudy (anonymous):

okay now it looks more readable

OpenStudy (anonymous):

now you can multiply by 4

OpenStudy (anonymous):

15i*4=60i right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

okay thank you for the help so much

OpenStudy (anonymous):

no problem